\begin{thebibliography}{19}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\providecommand{\urlprefix}{URL }
\providecommand{\selectlanguage}[1]{\relax}

\bibitem[{Babai \emph{et~al.}(1986)Babai, Frankl \&
  Simon}]{DBLP:conf/focs/BabaiFS86}
\textsc{L{\'a}szl{\'o} Babai}, \textsc{Peter Frankl} \& \textsc{Janos Simon}
  (1986).
\newblock Complexity classes in communication complexity theory (preliminary
  version).
\newblock In \emph{27th Annual Symposium on Foundations of Computer Science,
  FOCS 1986, 27--29 October 1986, Toronto, Ontario, Canada}, 337--347. IEEE
  Computer Society.

\bibitem[{Balc{\'a}zar \emph{et~al.}(1990)Balc{\'a}zar, D{\'\i}az \&
  Gabarr{\'o}}]{books/SCII}
\textsc{Jos{\'e}~L. Balc{\'a}zar}, \textsc{Josep D{\'\i}az} \& \textsc{Joaquim
  Gabarr{\'o}} (1990).
\newblock \emph{Structural Complexity II}.
\newblock Texts in Theoretical Computer Science, An EATCS Series.
  Springer-Verlag, 1st edition.

\bibitem[{Balc{\'a}zar \emph{et~al.}(1995)Balc{\'a}zar, D{\'\i}az \&
  Gabarr{\'o}}]{books/SCI}
\textsc{Jos{\'e}~L. Balc{\'a}zar}, \textsc{Josep D{\'\i}az} \& \textsc{Joaquim
  Gabarr{\'o}} (1995).
\newblock \emph{Structural Complexity I}.
\newblock Texts in Theoretical Computer Science, An EATCS Series.
  Springer-Verlag, 2nd edition.

\bibitem[{Beigel \emph{et~al.}(1991)Beigel, Reingold \&
  Spielman}]{DBLP:conf/stoc/BeigelRS91}
\textsc{Richard Beigel}, \textsc{Nick Reingold} \& \textsc{Daniel~A. Spielman}
  (1991).
\newblock PP Is Closed Under Intersection (Extended Abstract).
\newblock In \emph{Proceedings of the Twenty Third Annual ACM Symposium on
  Theory of Computing, 6-8 May 1991, New Orleans, Louisiana, USA}, 1--9. ACM.

\bibitem[{Du \& Ko(2000)}]{books/KIKCC}
\textsc{Ding-Zhu Du} \& \textsc{Ker-I Ko} (2000).
\newblock \emph{Theory of Computational Complexity}.
\newblock Wiley-Interscience Series in Discrete Mathematics and Optimization.
  John Wiley \& Sons, Inc., 1st edition.

\bibitem[{Fenner \emph{et~al.}(1991)Fenner, Fortnow \&
  Kurtz}]{DBLP:conf/coco/FennerFK91}
\textsc{Stephen~A. Fenner}, \textsc{Lance Fortnow} \& \textsc{Stuart~A. Kurtz}
  (1991).
\newblock Gap-Definable Counting Classes.
\newblock In \emph{Structure in Complexity Theory Conference}, 30--42.

\bibitem[{Hemaspaandra \& Ogihara(2002)}]{books/ComplexityTheoryCompanion}
\textsc{Lane~A. Hemaspaandra} \& \textsc{Mitsunori Ogihara} (2002).
\newblock \emph{The Complexity Theory Companion}.
\newblock Texts in Theoretical Computer Science, An EATCS Series.
  Springer-Verlag.

\bibitem[{Klauck(2001)}]{DBLP:conf/focs/Klauck01}
\textsc{Hartmut Klauck} (2001).
\newblock Lower Bounds for Quantum Communication Complexity.
\newblock In \emph{42nd Annual Symposium on Foundations of Computer Science,
  October 14--17, Las Vegas Nevada, USA}, 288--297. IEEE.

\bibitem[{Klauck(2003)}]{DBLP:conf/coco/Klauck03}
\textsc{Hartmut Klauck} (2003).
\newblock Rectangle Size Bounds and Threshold Covers in Communication
  Complexity.
\newblock In \emph{18th Annual IEEE Conference on Computational
  Complexity,7--10 July 2003, Aarhus, Denmark}, 118--134. IEEE Computer
  Society.

\bibitem[{K{\"o}bler \emph{et~al.}(1993)K{\"o}bler, Sch{\"o}ning \&
  Tor{\'a}n}]{books/bh/KSTGraphIso}
\textsc{Johannes K{\"o}bler}, \textsc{Uwe Sch{\"o}ning} \& \textsc{Jacobo
  Tor{\'a}n} (1993).
\newblock \emph{The Graph Isomorphism Problem -- Its Structural Complexity}.
\newblock Birkh{\"a}user Boston.

\bibitem[{Kushilevitz \& Nisan(1997)}]{books/KushilevitzNisan}
\textsc{Eyal Kushilevitz} \& \textsc{Noam Nisan} (1997).
\newblock \emph{Communication Complexity}.
\newblock Cambridge University Press.

\bibitem[{Linial \& Shraibman(2009)}]{DBLP:journals/cpc/LinialS09}
\textsc{Nathan Linial} \& \textsc{Adi Shraibman} (2009).
\newblock Learning Complexity vs Communication Complexity.
\newblock \emph{Combinatorics, Probability {\&} Computing} \textbf{18}(1--2),
  227--245.

\bibitem[{Lokam(1995)}]{DBLP:conf/focs/Lokam95}
\textsc{Satyanarayana~V. Lokam} (1995).
\newblock Spectral Methods for Matrix Rigidity with Applications to Size-Depth
  Tradeoffs and Communication Complexity.
\newblock In \emph{36th Annual Symposium on Foundations of Computer Science,
  Milwaukee, 23--25 October 1995}, 6--15. IEEE Computer Society.

\bibitem[{Lokam(2001)}]{DBLP:journals/jcss/Lokam01}
\textsc{Satyanarayana~V. Lokam} (2001).
\newblock Spectral Methods for Matrix Rigidity with Applications to Size-Depth
  Trade-offs and Communication Complexity.
\newblock \emph{J. Comput. Syst. Sci.} \textbf{63}(3), 449--473.

\bibitem[{Razborov(1989)}]{RazborovRigidMatrices}
\textsc{Alexander Razborov} (1989).
\newblock {On Rigid Matrices (in Russian)}.
\newblock Technical report, Steklov Mathematical Institute.
\newblock Electronically available at
  \url{http://people.cs.uchicago.edu/~razborov/rigid.pdf}.

\bibitem[{Tarui(1991)}]{DBLP:conf/stacs/Tarui91}
\textsc{Jun Tarui} (1991).
\newblock Randomized Polynomials, Threshold Circuits, and the Polynomial
  Hierarchy.
\newblock In \emph{STACS 91, 8th Annual Symposium on Theoretical Aspects of
  Computer Science, Hamburg, Germany, February 14-16, 1991, Proceedings},
  \textsc{Christian Choffrut} \& \textsc{Matthias Jantzen}, editors, volume 480
  of \emph{Lecture Notes in Computer Science}, 238--250. Springer.

\bibitem[{Toda(1991)}]{DBLP:journals/siamcomp/Toda91}
\textsc{Seinosuke Toda} (1991).
\newblock {PP} is as Hard as the Polynomial-Time Hierarchy.
\newblock \emph{SIAM J. Comput.} \textbf{20}(5), 865--877.

\bibitem[{Wunderlich(2010)}]{DBLP:journals/eccc/Wunderlich10}
\textsc{Henning Wunderlich} (2010).
\newblock On a Theorem of Razborov.
\newblock \emph{Electronic Colloquium on Computational Complexity (ECCC)}
  \textbf{17}, 86.

\bibitem[{Yao(1979)}]{DBLP:conf/stoc/Yao79}
\textsc{Andrew Chi-Chih Yao} (1979).
\newblock Some Complexity Questions Related to Distributive Computing
  (Preliminary Report).
\newblock In \emph{Conference Record of the Eleventh Annual ACM Symposium on
  Theory of Computing, 30 April--2 May, 1979, Atlanta, Georgia, USA}, 209--213.
  ACM.

\end{thebibliography}

\documentclass[a4paper,10pt,twoside,onecolumn,notitlepage,final]{article}

\usepackage[latin1]{inputenc}
\usepackage[T1]{fontenc}
\usepackage[english]{babel}

\usepackage[noccalgorithms]{cc-cls}
\usepackage{amsmath} 
\usepackage{xspace}
\usepackage{mathtools}

\providecommand{\mrm}[1]{\ensuremath{\mathrm{#1}}}

\providecommand{\mcal}[1]{\ensuremath{\mathcal{#1}}} 
\providecommand{\mbb}[1]{\ensuremath{\mathbb{#1}}}

\DeclareMathOperator{\lc}{lc}

\newcommand{\opBP}{\mcal{BP}\cdot}

\newcommand{\ccc}{\ensuremath{\mrm{cc}}}

\newcommand{\sPcc}{\#\mcal{P}^{\ccc}}

\newcommand{\PPcc}{\mcal{PP}^{\ccc}}
\newcommand{\PHcc}{\mcal{PH}^{\ccc}}
\newcommand{\PSPACEcc}{\mcal{PSP\!ACE}^{\ccc}}

\newcommand{\PP}{\mrm{PP}}
\newcommand{\disc}{\mrm{disc}}
\newcommand{\mc}{\mrm{mc}}

\newcommand{\BP}{\mrm{BP}}

\newcommand{\acc}{\mrm{acc}}
\newcommand{\rej}{\mrm{rej}}
\newcommand{\gap}{\mrm{gap}}

\newcommand{\ABie}{i.e.,\ } 
\newcommand{\ABeg}{e.g.,\ }
\newcommand{\ABvs}{vs.\xspace}
\newcommand{\ABp}{p.\xspace}
\newcommand{\ABWlog}{W.l.o.g.,\ }
\newcommand{\ABwlog}{w.l.o.g.\ }

\newcommand{\B}{\mbb{B}}

\newcommand{\X}{\mcal{X}}
\newcommand{\Y}{\mcal{Y}}

\newcommand{\XtY}{\X\times\Y}

\newcommand{\Mid}{\,\middle|\,}

\newcommand{\das}{\vcentcolon\nolinebreak\mkern-1.2mu\nolinebreak=}

\newcommand{\polylog}{\mrm{polylog}}

\title{A note on a problem in communication complexity}

\author{Henning Wunderlich\footnote{Capgemini, Carl-Wery-Str. 42, 81739 München, Email: \texttt{henning.wunderlich@capgemini.com}}}

\begin{document}
\maketitle
\begin{abstract}
\noindent
In this note, we prove a version of Tarui's Theorem in communication complexity, namely $\PHcc \subseteq \opBP\PPcc$.
Consequently, every measure for $\PPcc$ leads to a measure for $\PHcc$, subsuming a result of Linial
and Shraibman that problems with high mc-rigidity lie outside the polynomial hierarchy.
By slightly changing the definition of mc-rigidity (arbitrary instead of uniform distribution), it is then evident
that the class $\mcal{M}^{\ccc}$ of problems with low mc-rigidity equals $\opBP\PPcc$. As
$\opBP\PPcc \subseteq \PSPACEcc$, this rules out the possibility, that had been left open, that even polynomial space
is contained in $\mcal{M}^{\ccc}$.
\end{abstract}

\section{Introduction}
This note is a contribution to the field of communication complexity.
We refer the reader to \cite{books/KushilevitzNisan} for an excellent introduction.
We are concerned with ideas circling around the $\PHcc$-\ABvs-$\PSPACEcc$ problem,
a long-standing open problem in structural communication complexity, first posed in \cite{DBLP:conf/focs/BabaiFS86}.

For each computation model there exists a corresponding structural complexity theory.
The study of structural complexity theory began by considering circuit classes and the Turing machine model,
see \ABeg \cite{books/ComplexityTheoryCompanion,books/KIKCC,books/SCI,books/SCII}
for good introductions. A prominent result in this area, influenced by \cite{DBLP:journals/siamcomp/Toda91},
is Tarui's Theorem, see \cite{DBLP:conf/stacs/Tarui91},
relating the polynomial hierarchy to probabilistic computation modes.

Starting with \cite{DBLP:conf/focs/BabaiFS86}, communication complexity classes
were defined and their relationships were studied. In contrast to the Turing-machine model, much is known
about the relationships between the set of standard classes for the communication model (Yao's model, \citealp{DBLP:conf/stoc/Yao79}). 
Unfortunately, the difficulties start with the second level of
the polynomial hierarchy and, as said before, it is a long-standing open problem,
whether or not the polynomial hierarchy, $\PHcc$, and polynomial space, $\PSPACEcc$, differ. 

Several strategies have been proposed to tackle this problem.
Razborov's strategy is based on the rigidity of finite-field rank
(see \citealp{RazborovRigidMatrices,DBLP:journals/eccc/Wunderlich10}).
Lokam uses ideas of Tarui to reduce the problem to rigidity problems, where those ridigities are defined via rank over the
field of real numbers (\citealp{DBLP:journals/jcss/Lokam01}).
\citealp{DBLP:journals/cpc/LinialS09} establish a connection to learning theory.
They define the notion of mc-rigidity and show that high mc-ridigity yields problems outside the polynomial hierarchy.
Furthermore, they conjecture that families of Hadamard matrices have high mc-rigidity. If true, this would yield the desired
separation.

In this note, we prove a version of Tarui's Theorem in communication complexity, namely $\PHcc \subseteq \opBP\PPcc$.
Consequently, every measure for $\PPcc$ leads to a measure for $\PHcc$, subsuming one of the results of Linial
and Shraibman mentioned above.
We slightly change the definition of mc-rigidity. In our terminology, we apply the BP-operator on margin complexity.
Now, an arbitrary probability distribution is allowed in the definition of margin rigidity.
In contrast, in the original definition, the uniform distribution was used.
(Hence, it is possible to consider unbalanced communication matrices as candidates for high mc-rigidity, too.)
It is then evident that the class $\mcal{M}^{\ccc}$ of problems with low mc-rigidity equals $\opBP\PPcc$. As
$\opBP\PPcc \subseteq \PSPACEcc$, this rules out the possibility, that had been left open by prior work, that even polynomial space
is contained in $\mcal{M}^{\ccc}$. In other words, the possibility that mc-rigidity defines
a communication complexity class, which is too big to be useful for the desired separation result between
$\PHcc$ and $\PSPACEcc$, is ruled out.

\section{Structural complexity theory}
\subsection{On $\PPcc$}
In the setting of communication complexity, formal languages are defined a bit differently than in the Turing-machine
world. Let $\B \das \{0,1\}$ denote the Boolean alphabet.
The set of pairs of strings of equal length is denoted by
$\B^{\ast\ast} \das \{ (x,y) \mid x,y \in \B^{\ast}, |x| = |y|\}$.
A \emph{(formal) language} $L$ is a subset of $\B^{\ast\ast}$, its \emph{$n$-bit section}
$L_{n}$ is the set of all pairs $(x,y) \in L$ of $n$-bit words $x,y$.
A \emph{communication complexity class} is a set of languages.

Communication complexity classes were first defined in \cite{DBLP:conf/focs/BabaiFS86}, in particular, the analog of probabilistic polynomial
time, $\PPcc$. In this subsection, we recall basic definitions and properties related to this class.

We define a \emph{guess protocol} $\Pi$ (over domain $\XtY$ with range $\B$) as a finite sequence $\Pi \das (\Pi_{1},\ldots,\Pi_{l})$ 
of deterministic protocols $\Pi_{i}$ (over domain $\XtY$ with range $\B$). We say that $\Pi$ \emph{uses} $l$ guesses.

The \emph{number of accepting guesses} of $\Pi$ on input $(x,y)$ is defined as
\begin{equation*}
\acc_{\Pi}(x,y) \das \left|\left\{ i \in [l] \Mid f_{\Pi_{i}}(x,y) = 1 \right\}\right| = \sum_{i \in [l]} f_{\Pi_{i}}(x,y)
\enspace,
\end{equation*}
where $f_{\Pi_{i}}$ denotes the function computed by the deterministic protocol $\Pi_{i}$.

The \emph{number of rejecting guesses}, $\rej_{\Pi}(x,y)$, is defined analogously. Clearly, we have
$\acc_{\Pi}(x,y) + \rej_{\Pi}(x,y) = l$.

An \emph{acceptance mode} is a two-ary predicate. The only acceptance mode of interest in this work is the \emph{PP acceptance mode},
$\PP(\acc, \rej) \das (\acc > \rej)$.

A guess protocol $\Pi$ \emph{computes} a Boolean function $f$ in acceptance mode $\Xi$, if
\begin{equation*}
f(x,y) = 1 \iff \Xi\left(\acc_{\Pi}(x,y), \rej_{\Pi}(x,y)\right)
\enspace.
\end{equation*}
Given a guess protocol $\Pi$ and an acceptance mode $\Xi$, we denote by $f^{\Xi}_{\Pi}$ the Boolean function computed
by $\Pi$ in acceptance mode $\Xi$.

The \emph{(worst-case) communication cost}, $\PP(\Pi)$, of a guess protocol $\Pi$ is defined as
$\PP(\Pi) \das \lceil \log l \rceil + \max_{i \in [l]}\mrm{D}(\Pi_{i})$, where $\mrm{D}(\Pi_{i})$ denotes the worst-case communication cost of the
deterministic protocol $\Pi_{i}$.

The \emph{(worst-case) PP communication complexity}, $\PP(f)$, of a Boolean function $f$ is defined as the minimum worst-case
communication cost of a guess protocol computing $f$ in PP acceptance mode.

A family of guess protocols $\Pi = (\Pi_{n})_{n \geq 1}$ is called \emph{efficient},
if the communication cost of $\Pi_{n}$ is $\polylog(n)$.

The set $\sPcc$ is defined as the set of function families $f = (f_{n})_{n \geq 1}$ such that there exists an efficient family
of guess protocols $\Pi = (\Pi_{n})_{n \geq 1}$ with $f_{n} = \acc_{\Pi_{n}}$.

The communication complexity class $\PPcc$ is then defined as the set of languages with efficient PP communication complexity, \ABie
\begin{equation*}
\PPcc \das \{ L \mid \PP(L_{n}) = \polylog(n) \}
\enspace.
\end{equation*}
  
The following result is well-known in structural complexity theory. We give a proof for the sake of completeness.

\begin{proposition}\label{Prop:PP:Equivalence}
For a language $L$ the following statements are equivalent.
\begin{enumerate}
\item $L \in \PPcc$.
\item $\exists f = (f_{n})_{n \geq 1} \in \sPcc\colon\exists g = (g_{n})_{n \geq 1}\colon g_{n} \leq 2^{\polylog(n)}$
such that for all $n$-bit inputs $(x,y)$ we have
\begin{equation*}
(x,y) \in L_{n} \iff f_{n}(x,y) > g_{n}
\enspace.
\end{equation*}
\end{enumerate}
\end{proposition}

\begin{proof}
\fbox{$\Rightarrow$} Let $\Pi \das (\Pi_{n})_{n \geq 1}$ be an efficient family of guess protocols
$\Pi_{n} \das (\Pi^{(n)}_{1},\ldots,\Pi^{(n)}_{l_{n}})$ computing $L$ in PP acceptance mode. Efficiency implies $l_{n} \leq 2^{\polylog(n)}$.
Define $f_{n} \das \acc_{\Pi_{n}}$ and $g_{n} \das \lfloor l_{n} / 2 \rfloor$. Then $f \das (f_{n})_{n \geq 1} \in \sPcc$ and
\begin{align*} 
(x,y) \in L_{n} &\iff \acc_{\Pi_{n}}(x,y) > \rej_{\Pi_{n}}(x,y)\\
&\iff \acc_{\Pi_{n}}(x,y) > l_{n} - \acc_{\Pi_{n}}(x,y)\\
&\iff f_{n}(x,y) > g_{n}
\enspace.
\end{align*}
\fbox{$\Leftarrow$} Let $f = (f_{n})_{n \geq 1} \in \sPcc$ and $g = (g_{n})_{n \geq 1}$, $g_{n} \leq 2^{\polylog(n)}$ be given.
Then there exists an efficient family $\Pi = (\Pi_{n})_{n \geq 1}$ of guess protocols
$\Pi_{n} \das (\Pi^{(n)}_{1},\ldots,\Pi^{(n)}_{l_{n}})$ such that $f_{n} = \acc_{\Pi_{n}}$. \ABWlog we can assume that
$l_{n} \geq 2g_{n}$. Otherwise, we add $(2g_{n} - l_{n})$ many trivial, always-rejecting protocols to $\Pi_{n}$.
We define an efficient family $\widetilde{\Pi} \das (\widetilde{\Pi}_{n})_{n \geq 1}$ of guess protocols as follows.
The protocol $\widetilde{\Pi}_{n}$ consists of the sequence of deterministic protocols in $\Pi_{n}$ plus $(l_{n} - 2g_{n})$
many trivial, always-accepting protocols. Clearly, we have
\begin{align*}
\acc_{\widetilde{\Pi}_{n}}(x,y) &= \acc_{\Pi_{n}}(x,y) + l_{n} - 2g_{n}
\enspace,\\
\rej_{\widetilde{\Pi}_{n}}(x,y) &= \rej_{\Pi_{n}}(x,y) = l_{n} - \acc_{\Pi_{n}}(x,y)
\enspace.
\end{align*}
As a consequence, we obtain
\begin{align*}
\acc_{\widetilde{\Pi}_{n}}(x,y) > \rej_{\widetilde{\Pi}_{n}}(x,y)
&\iff \acc_{\Pi_{n}}(x,y) + l_{n} - 2g_{n} > l_{n} - \acc_{\Pi_{n}}(x,y)\\
&\iff \acc_{\Pi_{n}}(x,y) > g_{n}\\
&\iff (x,y) \in L_{n}
\enspace,
\end{align*}
where the last equivalence is by the assumption. Hence, $\widetilde{\Pi}$ is an efficient family of guess protocols
computing $L$ in PP acceptance mode, \ABie $L \in \PPcc$. 
\end{proof}

In the remaining part of this subsection, we transfer results of \cite{DBLP:conf/stoc/BeigelRS91} to communication complexity.

Given a guess protocol $\Pi \das (\Pi_{1},\ldots,\Pi_{l})$, we consider the guess protocol $\Pi' \das (\Pi_{1},\Pi_{1},\ldots,\Pi_{l},$ $\Pi_{l},0)$,
where $0$ denotes the always-rejecting protocol. Then $\Pi'$ has the property that $\acc_{\Pi'}(x,y) \neq \rej_{\Pi'}(x,y)$
for all input pairs $(x,y)$, and $\Pi'$ computes the same function as $\Pi$ in PP acceptance mode.
Hence, \ABwlog we can assume that every guess protocol has the above property.

We adapt the convenient notation of \cite{DBLP:conf/coco/FennerFK91} and define $\gap_{\Pi}(x,y) \das \acc_{\Pi}(x,y) - \rej_{\Pi}(x,y)$.
Then, we have
\begin{align*}
f^{\PP}_{\Pi}(x,y) = 1 &\implies \gap_{\Pi}(x,y) > 0
\enspace,\\
f^{\PP}_{\Pi}(x,y) = 0 &\implies \gap_{\Pi}(x,y) < 0
\enspace.   
\end{align*}

For a deterministic protocol $\Pi$ with range $\B$, we define its \emph{complement} $\overline{\Pi}$ as the protocol which accepts
iff $\Pi$ rejects. Given a guess protocol $\Pi \das (\Pi_{1},\ldots,\Pi_{l})$, we define its \emph{complement}, $\overline{\Pi}$, as
$\overline{\Pi} \das (\overline{\Pi}_{1},\ldots,\overline{\Pi}_{l})$. Clearly, we have $\gap_{\overline{\Pi}} = -\gap_{\Pi}$.

Given two guess protocols $\Pi \das (\Pi_{1},\ldots,\Pi_{l_1})$ and $\Pi' \das (\Pi'_{1},\ldots,\Pi'_{l_2})$, respectively, we define their
\emph{sum}, $\Pi + \Pi'$, as $\Pi + \Pi' \das (\Pi_{1},\ldots,\Pi_{l_1},\Pi'_{1},\ldots,\Pi'_{l_2})$. Here, we have
$\gap_{\Pi + \Pi'} = \gap_{\Pi} + \gap_{\Pi'}$.

Let $\Pi$ and $\Pi'$ be two deterministic protocols. We define their \emph{product}, $\Pi * \Pi'$, as the deterministic protocol, which runs
as follows. First, $\Pi$ is executed. If $\Pi$ accepts, then $\Pi'$ is executed, else $\overline{\Pi'}$ is executed.
Given two guess protocols $\Pi \das (\Pi_{1},\ldots,\Pi_{l_1})$ and $\Pi' \das (\Pi'_{1},\ldots,\Pi'_{l_2})$, respectively, we define their
\emph{product}, $\Pi * \Pi'$, as $\Pi * \Pi' \das (\Pi_{1} * \Pi'_{1},\ldots,\Pi_{1} * \Pi'_{l_2},\ldots,\Pi_{l_1} * \Pi'_{l_2})$.
In this case, $\gap_{\Pi * \Pi'} = \gap_{\Pi} \cdot \gap_{\Pi'}$.

The following lemma corresponds to \cite[Lemma 5]{DBLP:conf/stoc/BeigelRS91}.

\begin{lemma}\label{PP:Lemma1}
Let $\Pi_{1},\ldots,\Pi_{k}$ be guess protocols using at most $l$ guesses and having communication cost at most $c$.
Let $p(z_{1},\ldots,z_{k})$ be a polynomial of degree $d$ with integer coefficients bounded above in absolute value by $M$.
Then there exists a guess protocol $\Pi$ such that
\begin{equation*}
\gap_{\Pi}(x,y) = p\left(\gap_{\Pi_{1}}(x,y),\ldots,\gap_{\Pi_{k}}(x,y)\right)
\enspace,
\end{equation*}
$\Pi$ uses at most
\begin{equation*}
M l^{d}(d + k)^{k + 1}
\end{equation*}
many guesses, and has communication cost bounded above by
\begin{equation*}
\left\lceil \log M + d\log l + (k+1)\log(d + k) \right\rceil + cd
\enspace.
\end{equation*}
\end{lemma}

\begin{proof}
The guess protocol $\Pi$ first guesses a monomial of $p$, say $\pm c z^{\alpha_{1}}_{1} \cdots z^{\alpha_{k}}_{k}$, with $c > 0$.
As $p$ has at most $\sum_{i = 0}^{d}{i + k \choose k}$ monomials, this requires using $\leq (d + k)^{k + 1}$ guesses.
Then $\Pi$ guesses one of $c$ branches, computes the product as described above, and complements if necessary.
Here $\Pi$ uses $M \cdot l^{d}$ additional guesses.
\end{proof}

The \emph{degree} of a rational function is defined as the maximum of the degrees of its numerator and denominator.
The following lemma corresponds to \cite[Lemma 6]{DBLP:conf/stoc/BeigelRS91}.

\begin{lemma}\label{PP:Lemma2}
Let $\Pi_{1},\ldots,\Pi_{k}$ be guess protocols using at most $l$ guesses and having communication cost at most $c$.
Let $r(z_{1},\ldots,z_{k})$ a rational function of degree $d$ with integer coefficients bounded above in absolute value by $M$.
Then there exists a guess protocol $\Pi$ such that $\gap_{\Pi}(x,y)$ and
\begin{equation*}
r\left(\gap_{\Pi_{1}}(x,y),\ldots,\gap_{\Pi_{k}}(x,y)\right)
\end{equation*}
have the same sign for all $(x,y)$ where the latter is defined, $\Pi$ uses at most
\begin{equation*}
\left( M l^{d} (2d + k)^{k + 1} \right)^{2}
\end{equation*}
many guesses, and has communication cost bounded above by
\begin{equation*}
2 \left( \left\lceil \log M + d\log l + (k+1)\log(2d + k) \right\rceil + cd \right)
\enspace.
\end{equation*}
\end{lemma}

\begin{proof}
Let $r = p / q$.
We just apply \ref{PP:Lemma1} with polynomial $p \cdot q$. The degree of this polynomial is at most $2d$ and the absolute values of its
coefficients are bounded above by $M^{2} \cdot {2d + k \choose k}$.
\end{proof}

The following functions are defined and studied in \cite{DBLP:conf/stoc/BeigelRS91}.
\begin{align*}
P_{m}(z) &\das (z - 1)\prod_{i = 1}^{m}(z - 2^{i})^{2}
\enspace,\\
S^{(k)}_{m}(z) &\das \frac{\left(P_{m}(-z)\right)^{h(k)} - \left(P_{m}(z)\right)^{h(k)}}
{\left(P_{m}(-z)\right)^{h(k)} + \left(P_{m}(z)\right)^{h(k)}}
\enspace,\\
T^{(k)}_{m}(z_{1},\ldots,z_{k}) &\das 2S^{(2k)}_{m}(z_{1}) + \cdots + 2S^{(2k)}_{m}(z_{k}) + 1
\enspace.
\end{align*}
Here, $h(k)$ denotes the least odd integer greater than or equal to $\log(2k + 1)$.

The following proposition corresponds to \cite[Lemma 9 and 10]{DBLP:conf/stoc/BeigelRS91}.
 
\begin{proposition}\mbox{}\label{Prop:DegM}
\begin{enumerate}
\item The degree of $P^{h(k)}_{m}$ is $h(k)(2m + 1)$ and the absolute value of each of its coefficients is bounded by
$2^{2h(k)\log h(k)) + 3h(k)m\log(2m+1)}$.
\item If $1 \leq z \leq 2^{m}$ then $1 \leq S^{(k)}_{m}(z) < 1 + 1/k$. If $-2^{m} \leq z \leq -1$ then $-1 - 1/k < S^{(k)}_{m}(z) \leq -1$.
The rational function $S^{(k)}_{m}(z)$ has degree $\leq h(k)(2m + 1)$ and the absolute value of each of its coefficients is bounded by
$2^{1 + 2h(k)\log h(k)) + 3h(k)m\log(2m+1)}$.
\item Assume that $1 \leq |z_{i}| \leq 2^{m}$ for $1 \leq i \leq k$. Then $T^{(k)}_{m}$ is a rational function that is positive
if at last half of the $z_{i}$'s are positive, and negative otherwise. The degree of $T^{(k)}_{m}$ is $\leq h(2k)(2m + 1)$, and the absolute
value of each of its coefficients is bounded by $2^{3h(2k)(\log h(2k) + m\log(2m+1) + 1)}$.
\end{enumerate}
\end{proposition}

A proof is given in the appendix.

\begin{proposition}\label{Prop:Majority}
Let $\Pi_{1},\ldots,\Pi_{k}$ be PP protocols with communication cost at most $c$. Then there exists a PP protocol $\Pi$ with
communication cost at most $\mcal{O}\!\left(k(\log k)(\log c) + c^{2}(\log k)\right)$ such that $\gap_{\Pi}(x,y)$ and
$T^{(k)}_{c}(\gap_{\Pi_{1}}(x,y),\ldots,\gap_{\Pi_{k}}(x,y))$ have the same sign for all input pairs $(x,y)$.
\end{proposition}

\begin{proof}
We apply \ref{PP:Lemma2} on $\Pi_{1},\ldots,\Pi_{k}$ and $r \das T^{(k)}_{c}$ of degree $d \das \deg r \leq h(2k)(2c+1)$, where the absolute value of each
of the coefficients in $r$ is bounded by
\begin{equation*}
M \leq 2^{3h(2k)(\log h(2k) + c\log(2c+1) + 1)}
\enspace.
\end{equation*}
The protocols use at most $c$ guesses.
We obtain the desired protocol $\Pi$ with communication cost bounded above by
\begin{gather*}
2 \left( \left\lceil \log M + d\log l + (k+1)\log(2d + k) \right\rceil + cd \right)\\
\leq 2\left(3h(2k)(\log h(2k) + c\log(2c+1) + 1) + h(2k)(2c+1)(\log c) \right.\\
\left.{}+ (k+1)\log(2h(2k)(c+1) + k) + h(2k)c(2c+1) + 1 \right)\\
= \mcal{O}\!\left(k(\log k)(\log c) + c^{2}(\log k)\right)
\enspace.
\end{gather*} 
\end{proof}

This lays the ground for the probability amplification result for randomized PP-protocols stated in the next section.
 
\subsection{On $\opBP\PPcc$}
Assume we are given a computation model $C$ together with a cost function measuring the resources consumed during a computation.
First of all, this gives us a complexity measure $D$ by taking the infimum of the cost function over all ``$C$-machines'' in the computation model.
In addition, if we are given a notion of efficiency, we can define a complexity class $\mcal{C}$ including all decision problems $L$
with a complexity $D(L)$ that is considered efficient.
It is interesting to study the power of randomization by enriching the computation model $C$ with random bits. This can be done by
defining a \emph{random} $C$-machine as a probability distribution over $C$-machines, together with an acceptance mode, \ABeg bounded error.
Again, we have a cost measure, we can define a complexity measure $R$, and thus, we can also define a complexity class $\widetilde{\mcal{C}}$
including all decision problems $L$ with a complexity $R(L)$ that is considered efficient.
In structural complexity theory it has proven useful to define complexity class operators, \ABeg the BP-operator. In our case, this
operator describes the relationship between the complexity classes $\mcal{C}$ and $\widetilde{\mcal{C}}$, namely
$\widetilde{\mcal{C}} = \opBP\mcal{C}$. In communication complexity, we can even go one step further. Here, it is possible to
express the complexity measure $R$ as a perturbation of $D$, \ABie $R = \mrm{BP}\cdot D$.
Hence, we arrive at an equation like  
$\opBP\mcal{C} = \{ L \mid (\mrm{BP}\cdot D)(L) \text{ is efficient} \}$.
In the following, we work out all the details to obtain a precise statement of this kind for the class $\opBP\PPcc$.
 
A \emph{randomized} PP-protocol $\Pi$ (over domain $\XtY$ with range $\B$) is defined as a probability distribution $\alpha\colon A \to [0,1]$
over a finite set $\{ \Pi_{a} \mid a \in A \}$ of PP-protocols $\Pi_{a}$ (each over domain $\XtY$ with range $\B$).

We say that $\Pi$ \emph{computes} a Boolean function $f$ \emph{with (two-sided) $\epsilon$-error}, if for all $n$-bit input pairs $(x,y)$ we have
\begin{equation*}
\Pr_{\alpha}\left[ f^{\PP}_{\Pi_{\alpha}}(x,y) \neq f(x,y) \right] \leq \epsilon
\enspace.
\end{equation*}

The \emph{(worst-case) communication cost} of a randomized PP-protocol is defined as the maximum worst-case communication cost over all
PP-protocols with non-zero probability, \ABie $\mrm{BP\text{-}PP}(\Pi) \das \max\{ \PP(\Pi_{a}) \mid a \in A, \alpha(a) > 0 \}$.

The \emph{(worst-case) $\epsilon$-error BP-PP communication complexity}, $\mrm{BP\text{-}PP}_{\epsilon}(f)$, of a Boolean function $f$
is defined as the minimum worst-case communication cost
of a randomized PP-protocol computing $f$ with two-sided $\epsilon$-error. If $\epsilon$ is not mentioned, we assume $\epsilon = 1/3$.

Probability amplification is possible for randomzied PP-protocols. As a prerequisite for a proof of this, we need the following
Chernoff-like result, which can be found in \cite[\ABp 70, Lemma 2.14]{books/bh/KSTGraphIso}.
 
\begin{fact}\label{Fact:ProbabilityAmplification}
Let $E$ be an event that occurs with probability $\frac{1}{2} + \epsilon$,
$0 < \epsilon \leq \frac{1}{2}$. Then $E$ occurs within $t$ independent trials ($t$ odd)
at least $t/2$ times with probability at least $1 - \frac{1}{2}\cdot \left(1 - 4\cdot \epsilon^{2}\right)^{t/2}$.
\end{fact}

\begin{theorem}[Probability amplification]\label{Thm:PA}
For every Boolean function $f$ and every $\epsilon \in [0,1/2[$ we have
\begin{equation*}
\mrm{BP\text{-}PP}_{\frac{1}{2}\cdot \left(1 - 4\cdot \epsilon^{2}\right)^{k/2}}(f)
\leq
\mcal{O}\left(k(\log k)\mrm{BP\text{-}PP}^{2}_{\epsilon}(f)\right)
\enspace.
\end{equation*}
\end{theorem}

\begin{proof}
Apply \ref{Prop:Majority} and \ref{Fact:ProbabilityAmplification}.
\end{proof}
 
We recall the definition of the BP-operator $\opBP\mcal{C}$ for communication complexity classes $\mcal{C}$ given in
\cite{DBLP:journals/eccc/Wunderlich10}.
 
A language $L$ is in $\opBP\mcal{C}$ if there exist a language $L' \in \mcal{C}$ and a polynomially bounded function  $q$
such that for all $n$-bit input pairs $(x,y)$ we have
\begin{align*}
(x,y) \notin L &\implies \left|\left\{ r \in \B^{\lceil q(\log n)\rceil} \Mid
(\langle x,r\rangle,\langle y,r\rangle) \in L' \right\}\right| \big/
2^{\lceil q(\log n)\rceil } \leq 1 / 3
\enspace,\\
(x,y) \in L &\implies
\left|\left\{ r \in \B^{\lceil q(\log n)\rceil} \Mid
(\langle x,r\rangle,\langle y,r\rangle) \in L' \right\}\right| \big/
2^{\lceil q(\log n)\rceil} \geq 2 / 3
\enspace.
\end{align*}

\begin{claim}\label{Claim:1}
\begin{equation*}
\opBP\PPcc = \{ L \mid \mrm{BP\text{-}PP}(L_{n}) = \mrm{polylog}(n) \}
\enspace.
\end{equation*}
\end{claim}

\begin{proof}
The proof of the $\subseteq$-inclusion is trivial. The other inclusion is an application of a result of Newman, see \ABeg
\cite[Theorem 3.14]{books/KushilevitzNisan}, which allows us to replace the arbitrarily large set $A$ and the distribution $\alpha$ of a
randomized PP-protocol by a uniform distribution on $\polylog(n)$ bits. We have to pay for this by increasing the error slightly,
but this is not a real problem any longer, because by \ref{Thm:PA} probability amplification is possible to reduce the error to less than one-third again. 
\end{proof}

Let $\Lambda\colon \mbb{M}_{n}(\{0,1\}) \to \mbb{R}$ be a mapping, assigning to each Boolean matrix $f$ a real number $\Lambda(f)$.
In the sequel, it will be the communication-complexity measure $\PP$. Let $\epsilon \in [0,1/2[$.
The \emph{BP-operator} applied on $\Lambda$ is defined as
\begin{equation*}
(\BP_{\epsilon}\cdot \Lambda)(f) \das \max_{\mu}\min_{\tilde{f}\colon \mu(f \neq \tilde{f}) \leq \epsilon}\Lambda(\tilde{f}) 
\enspace,
\end{equation*} 
where $\mu$ denotes a probability distribution on the matrix entries of $f$. Again, if $\epsilon$ is not mentioned, we assume
that $\epsilon = 1/3$.

We remark that the BP-operator may be considered as a perturbation operator that tests how much the measure $\Lambda$ deviates from
the value $\Lambda(f)$ when $f$ is altered by an $\epsilon$-fraction of its entries.
 
\begin{claim}\label{Claim:2}
For every Boolean function $f$ and every $\epsilon \in [0,1/2[$ we have
\begin{equation*}
\mrm{BP\text{-}PP}_{\epsilon}(f) = (\BP_{\epsilon}\cdot\PP)(f)
\enspace.
\end{equation*}
\end{claim}

\begin{proof}
This is just an application of Yao's Minimax-principle, see \ABeg \cite[Theorem 3.20]{books/KushilevitzNisan}.
Here, the PP-protocols take the role of the deterministic protocols in the original proof.
\end{proof}

Combining \bare\ref{Claim:1} and \bare\ref{Claim:2}, we obtain
 
\begin{proposition}
For every language $L$ we have
\begin{equation*}
L \in \opBP\PPcc \iff (\BP_{\epsilon}\cdot\PP)(L_{n}) = polylog(n)
\enspace,
\end{equation*}
where $\epsilon \in [0,1/2[$ is an arbitrary but fixed constant.
\end{proposition}

A result of Klauck gives a characterization of $\PP$-complexity via discrepancy, $\disc'(B)$, defined
for Boolean matrices $B$.

\begin{fact}[\cite{DBLP:conf/focs/Klauck01}, Fact 6 in \cite{DBLP:conf/coco/Klauck03}]\label{Fact:Klauck}
For every language $L$ we have
\begin{equation*}
\log \frac{1}{\disc'(L_{n})} \leq \PP(L_{n}) \leq \mcal{O}\left(\log\frac{1}{\disc'(L_{n})} + \log n\right)
\enspace. 
\end{equation*}
\end{fact}

A main result of \cite{DBLP:journals/cpc/LinialS09} is a tight relationship between margin complexity, $\mc(A)$,
and discrepancy, $\disc(A)$, defined for sign matrices $A$.

\begin{fact}{\cite[Theorem 3.4]{DBLP:journals/cpc/LinialS09}}\label{Fact:LS}
For a sign matrix $A$ the ratio between discrepany $\disc(A)^{-1}$ and margin complexity $\mc(A)$
is a factor of at most eight.
\end{fact}

The relationship between the Boolean and sign matrix version of discrepancy is given by
$\disc'(B) = \disc(J - 2B)$, where $J$ is the all-ones matrix.

Hence, combining \bare\ref{Fact:Klauck}, \bare\ref{Fact:LS} and \ref{TaruiTranslated},
we can rule out the possibility that polynomial space is strictly contained in the communication complexity class
implicitly defined by low mc-rigidity, here, defined as $(\BP\cdot\mc')(\cdot)$, where $\mc'(B) \das \mc(J - 2B)$
maps Boolean matrices to sign matrices.

\begin{corollary}
Let $\mcal{M}^{\ccc} \das \{ L \mid (\BP\cdot\mc')(L_{n}) = \mrm{polylog}(n)\}$. Then
$\mcal{M}^{\ccc} = \opBP\PPcc$.

Hence, the statement $\PSPACEcc \subsetneq \mcal{M}^{\ccc}$ is not true.
\end{corollary}

\subsection{On Tarui's Theorem}
A remarkable result in structural complexity theory is Toda's Theorem, which tells us that
the polynomial hierarchy $\mcal{PH}$ is contained in $\opBP\oplus\mcal{P}$, see \cite{DBLP:journals/siamcomp/Toda91}.
Using the concept of randomized polynomials, \cite{DBLP:conf/stacs/Tarui91} extended this further by showing
 
\begin{fact}[\cite{DBLP:conf/stacs/Tarui91}]
$\mcal{PH} \subseteq \opBP\mcal{PP}$.
\end{fact}

In fact, he even showed a stronger statement.

Often there are several pitfalls when one tries to transfer a result from structural complexity theory to communication complexity.
In case of Toda's Theorems, see \cite{DBLP:journals/eccc/Wunderlich10}, the use of complexity class operators was essential to avoid
problems with relativization. Establishing a communication complexity version of Tarui's result is a bit tricky, too.
Indeed, it is not at all clear how to transfer the proof of Theorem 4.1 in \cite{DBLP:conf/stacs/Tarui91} to Yao's model.
Instead, we express communication protocols as generalized $\mcal{AC}^{0}$ circuits and then apply Tarui's randomized
polynomial approximations for such circuits. As far as we know, the observation that languages in $\PHcc$ can be expressed by
$\mcal{AC}^{0}$ circuits is from \cite{RazborovRigidMatrices}. This was used by \cite{DBLP:conf/focs/Lokam95} together with
Tarui's result to prove upper bounds for weak rigidities for languages in $\PHcc$.
(Hence, lower bounds for weak rigidities would give us languages outside $\PHcc$.)
In the same vein, \cite{DBLP:journals/cpc/LinialS09} utilized this insight in the proof of their result that languages with
high mc-rigidity lie outside of the polynomial hierarchy.
Coming back again to randomized polynomial approximations, these objects are not $\sPcc$ functions yet. Hence, we have to apply
the same trick to handle the negative terms as in \cite[Theorem 3.2]{DBLP:conf/stacs/Tarui91}.

\begin{theorem}\label{TaruiTranslated}
$\PHcc \subseteq \opBP\PPcc$.
\end{theorem}

\begin{proof}
Let $L$ be a language in $\PHcc$. It was observed in \cite{RazborovRigidMatrices} and also \cite[Proof of Theorem 4.1]{DBLP:conf/focs/Lokam95}
that there exist an $\mcal{AC}^{0}$ circuit family $(C_{t_{n}})_{n \geq 1}$ of $\{ \vee , \wedge \}$ circuits $C_{t_{n}}$ of size $\leq 2^{\polylog(n)}$
with $t_{n} \leq 2^{\polylog(n)}$ many variables, and families $f^{(n)}_{1},\ldots,f^{(n)}_{t_{n}}$,
$g^{(n)}_{1},\ldots,g^{(n)}_{t_{n}}$ of Boolean functions such that for all $n$-bit input pairs $(x,y)$ we have
\begin{equation*}
(x,y) \in L_{n} \iff C_{t_{n}}\left(f^{(n)}_{1}(x) \wedge g^{(n)}_{1}(y),\ldots,f^{(n)}_{t_{n}}(x) \wedge g^{(n)}_{t_{n}}(y)\right) = 1
\enspace.
\end{equation*}
In \cite[Theorem 3.1]{DBLP:conf/stacs/Tarui91} it was shown that such a circuit family can be approximated with bounded error ($\epsilon = 1/3$)
by a family of randomized polynomials $\Phi_{t_{n}}$ over $\mbb{Z}$ such that the degree $d_{t_{n}}$ of $\Phi_{t_{n}}$ is
$d_{t_{n}} \leq \polylog(t_{n}) = \polylog(n)$, the absolute value of each coefficient is bounded by $2^{\polylog(t_{n})} = 2^{\polylog(n)}$,
$\Phi_{t_{n}}$ uses $\polylog(t_{n}) = \polylog(n)$ many random bits, and that $\Phi_{t_{n}}$ computes $C_{t_{n}}$ with two-sided error and Boolean guarantee.
We can write $\Phi_{n}(x,y) \das \Phi_{t_{n}}\left(f^{(n)}_{1}(x) \wedge g^{(n)}_{1}(y),\ldots,f^{(n)}_{t_{n}}(x) \wedge g^{(n)}_{t_{n}}(y)\right)$ as
\begin{align*}
\Phi_{n}(x,y) &= \sum_{S \subseteq [t_{n}], |S| \leq d_{t_{n}}}\alpha^{(n)}_{S} \cdot \prod_{s \in S}f^{(n)}_{s}(x)g^{(n)}_{s}(y)\\
&= \sum_{S \subseteq [t_{n}], |S| \leq d_{t_{n}}}\alpha^{(n)}_{S} \cdot f^{(n)}_{S}(x)g^{(n)}_{S}(y)
\enspace,
\end{align*}
where $f^{(n)}_{S}(x) \das \prod_{s \in S}f^{(n)}_{s}(x)$, and similarly for $g^{(n)}_{S}(y)$.

Define $g_{n}$ as the sum of the absolute values of the negative coefficients in $\Phi_{n}(x,y)$, and define
$\Psi_{n}(x,y) \das \Phi_{n}(x,y) + g_{n}$. Then
\begin{align*}
\Psi_{n}(x,y) &=
\sum_{\stackrel{S \subseteq [t_{n}], |S| \leq d_{t_{n}},}{\alpha^{(n)}_{S} > 0}}\ \sum_{1 \leq \alpha \leq \alpha^{(n)}_{S}}
f^{(n)}_{S}(x)g^{(n)}_{S}(y)\\
\\
&+
\sum_{\stackrel{S \subseteq [t_{n}], |S| \leq d_{t_{n}},}{\alpha^{(n)}_{S} < 0}}\ \sum_{1 \leq \alpha \leq \left|\alpha^{(n)}_{S}\right|}
\left(1 - f^{(n)}_{S}(x)g^{(n)}_{S}(y)\right)
\end{align*}
and $(\Psi_{n})_{n \geq 1}$ is clearly a family of randomized $\sPcc$ functions, because the number of terms is bounded by
\begin{equation*}
\max_{S}\left|\alpha^{(n)}_{S}\right| \cdot \left|t_{n} \choose d_{t_{n}}\right|
\leq 2^{\polylog(n)} \cdot 2^{\polylog(n) \cdot \polylog(n)}
= 2^{\polylog(n)}
\enspace.
\end{equation*}
In addition, for every $n$-bit input pair $(x,y)$ with high probability we have
\begin{equation*}
(x,y) \in L_{n} \iff \Psi_{n}(x,y) > g_{n}
\enspace.
\end{equation*}
Applying \ref{Prop:PP:Equivalence} on $(\Psi_{n})_{n \geq 1}$ and $(g_{n})_{n \geq 1}$
yields a randomized family of guess protocols computing $L$ in PP acceptance mode with bounded error. Hence, $L \in \opBP\PPcc$.
\end{proof}

\begin{corollary}
Any lower-bound method $\mu$ for $\PPcc$ leads to a lower-bound method $\mrm{BP}\cdot\mu$ for $\PHcc$ via perturbation with the BP-operator.  
\end{corollary}

In particular, this holds for instances such as the discrepancy method, the $\gamma^{\infty}_{2}$-norm and margin complexity,
the latter subsuming a result of Linial and Shraibman.

Let $\overline{L} \das \{ (x,y) \in \B^{\ast\ast} \mid (x,y) \notin L\}$ define the complement of a language $L$,
and let $\mrm{co}\cdot\mcal{C}$ denote the class of all complements of languages from $\mcal{C}$.

Define the \emph{RP-operator}, $\mcal{RP}\cdot\mcal{C}$, analogously to the BP-operator but with one-sided error.
Finally, define the \emph{Las-Vegas-operator}, $\mcal{ZP}\cdot$, as $\mcal{ZP}\cdot\mcal{C} \das \mcal{RP}\cdot\mcal{C} \cap \mrm{co}\cdot\mcal{RP}\cdot\mcal{C}$.

\begin{openquestion}
Do we have
$\PHcc \subseteq \mcal{ZP}\cdot\PPcc\ ?$
\end{openquestion}

\bibliography{note}

\appendix

\section{Proof of \ref{Prop:DegM}}
Denote by $\lc P(z)$ the absolute value of the largest coefficient of a univariate polynomial $P(z)$. Clearly, we have
\begin{equation*}
\lc P(z)Q(z) \leq (\deg P(z) + \deg Q(z))(\lc P(z))(\lc Q(z))
\enspace.
\end{equation*}
By induction, we obtain
\begin{equation*}
\lc P^{n}(z) \leq n!\left( \deg P(z) \cdot \lc P(z) \right)^{n}
\enspace.
\end{equation*}
\begin{enumerate}
\item We have $\deg P^{h(k)}_{m}(z) = h(k)\deg P_{m}(z) = h(k)(1 + m\cdot 2)$. Furthermore,
$\lc P_{m}(z) = 1 \cdot \lc \Pi_{i \in [m]}(z - 2^{i})^{2} \leq \lc (z - 2^{m})^{2m} \leq (2m)!(2m)2^{m}$.
Hence,
\begin{equation*}
\lc P^{h(k)}_{m}(z) \leq h(k)!\left(h(k)(2m+1) \cdot (2m)!(2m)2^{m}\right)^{h(k)} \leq
2^{2h(k)\log h(k) + 3h(k)m\log(2m+1)}
\enspace.
\end{equation*}

\item We have $\deg S^{(k)}_{m}(z) = \deg P^{h(k)}_{m}(z) = h(k)(2m+1)$ and
\begin{equation*}
\lc S^{(k)}_{m}(z) \leq 2\lc P^{h(k)}_{m}(z) \leq 2^{1 + 2h(k)\log h(k) + 3h(k)m\log(2m+1)}
\enspace.
\end{equation*}

\item Finally, we have $\deg T^{k}_{m}(z) = \deg S^{(2k)}_{m}(z) = h(2k)(2m+1)$ and
\begin{equation*}
\lc T^{(k)}_{m}(z) \leq 2\lc S^{(2k)}_{m}(z) \leq 2^{3h(k)(\log h(k) + m\log(2m+1) + 1)}
\enspace.
\end{equation*}
\end{enumerate}\qed

\end{document}